# Extendability of projective varieties via degeneration to ribbons with applications to Calabi-Yau threefolds

### v1

@@ 3. Calabi-Yau ribbons on Fano threefolds and deformations of Calabi-Yau double covers of Fano threefolds
Since $H^{2}(\mathcal{O}_{X})=0$, we have that the map $H^{1}(\mathcal{E}_{L})\to H^{1}(T_{X})$ between the first order deformations of the polarized pair $(X,lA)$ and the first order deformations of $X$ is surjective while the respective maps of obstruction spaces is injective. So we have that the forgetful map between the functor of deformations of the polarized pair to the functor of deformations of $X$ is smooth. By the $T^{1}-$ lifting criterion by [Ran92], $X$ is unobstructed and hence $(X,lA)$ is unobstructed as well. So the pair $(X,lA)$ is contained in a unique irreducible component of the moduli space of smooth polarized Calabi-Yau threefolds. By the condition of the $(4)$ both of the pairs $(X_{t},\mathcal{O}_{X_{t}}(1))$ and $(X^{\prime}_{s},\mathcal{O}_{X^{\prime}_{s}}(1))$ are also contained the same component of smooth polarized Calabi-Yau threefolds. So there is a one parameter family $(\mathcal{X},\mathcal{O}_{\mathcal{X}}(1))$ where $f:\mathcal{X}\to T$ is a scheme over $T$, such that $(\mathcal{X}_{t},\mathcal{O}_{\mathcal{X}_{t}}(1))=(X_{t},\mathcal{O}_{X_{t}}(1))$ while $(\mathcal{X}_{s},\mathcal{O}_{\mathcal{X}_{s}}(1))=(X^{\prime}_{s},\mathcal{O}_{X^{\prime}_{s}}(1))$. But now since $H^{i}(\mathcal{O}_{\mathcal{X}_{p}}(1))$ for $i>0$ for any $p\in T$, we have that $h^{0}(\mathcal{O}_{\mathcal{X}_{p}}(1))$ is constant for $p\in T$. Therefore $f_{*}(\mathcal{O}_{\mathcal{X}}(1))$ is locally free of rank $N_{l}+1$ and $\mathcal{X}\to\mathbb{P}(f_{*}(\mathcal{O}_{\mathcal{X}}(1)))$ is a $T-$ morphism for $T$ irreducible such that at $t\in T$ we get the embedding $X_{t}\subset\mathbb{P}^{N_{l}}$ while at $s\in T$, we get the embedding $X^{\prime}_{s}\subset\mathbb{P}^{N_{l}}$. Therefore the $X_{t}\subset\mathbb{P}^{N_{l}}$ and $X^{\prime}_{s}\subset\mathbb{P}^{N_{l}}$ belong to the same irreducible component of the Hilbert scheme. $\blacksquare$
@@ 4. Applications to extendability of Calabi-Yau threefolds and canonical surfaces
Proof. In the case, $1.2-1.4$, the general elements of $\mathcal{H}^{Y}_{1}$ are complete intersections and hence the result follows. We concentrate on $1.5-1.11$. We first show that in each case, a special locus in $\mathcal{H}^{Y}_{j}$ is $k-$ extendable for $k$ as mentioned in the table. In each of the cases the $k-$ th step of the respective embedding $Y\xhookrightarrow{|\mathcal{O}_{Y}(j)|}\mathbb{P}^{M}$ is a Fano variety $G\hookrightarrow\mathbb{P}^{M+k}$ as follows (see [M89A], [M89B] or [CLM98] or [Bel24]).
@@
This implies that $T_{\mathbb{P}^{M+k}|_{G}}(-1)$ is base point free and hence $N_{G/\mathbb{P}^{M+k}}(-1)$ is base point free. Further the rank of $N_{G/\mathbb{P}^{M}}(-1)$ which is $M-3$ is greater than $3$. Hence there exists a nowhere vanishing section of $H^{0}(N_{G/\mathbb{P}^{M+k}}(-1))$ which corresponds to an embedded ribbon $\widetilde{G}\subset\mathbb{P}^{M+k}$ on $G\subset\mathbb{P}^{M+k}$. Hence we get a very ample polarization $\mathcal{O}_{\widetilde{G}}(1)$ which restricts to $\mathcal{O}_{G}(1)$ on $G$. Note that $H^{0}(\mathcal{O}_{\widetilde{G}}(1))=H^{0}(\mathcal{O}_{G}(1))+1$ and hence the embedding $\widetilde{G}\subset\mathbb{P}^{M+k}$ is actually induced by a sublinear series of $\mathcal{O}_{\widetilde{G}}(1)$ of codimension $1$. Now embed $\widetilde{G}\subset\mathbb{P}^{M+k+1}$ by the complete linear series of $\mathcal{O}_{\widetilde{G}}(1)$. To show that $\widetilde{G}$ has an embedded smoothing in $\mathbb{P}^{M+k+1}$, we work as in Proposition 3.1. By [BGM24, Proposition $2.16$] (which uses [GGP13, Proposition $1.4$]) we need to show that $|\mathcal{O}_{G}(2)|$ has a smooth member and prove the vanishing of $H^{1}(\mathcal{O}_{G}(2)),H^{1}(N_{G/\mathbb{P}^{M}}(-1))$ and $H^{1}(N_{G/\mathbb{P}^{M}})$. The first one follows from Kodaira vanishing theorem. From the above two exact sequences, for the other two vanishings, one needs to show the vanishing of $H^{1}(T_{\mathbb{P}^{M}}|_{G}(-i))$ and $H^{2}(T_{G}(-i))$ for $i=0,1$. The first one boils down to the vanishings of $H^{1}(\mathcal{O}_{G}(i))$ for $i=0,1$ and $H^{2}(\mathcal{O}_{G}(-i))$ for $i=0,1$, both of which are true since $G$ is Fano. Finally $H^{2}(T_{G}(-i))=H^{d-2}(\Omega_{G}\otimes K_{G}\otimes\mathcal{O}_{G}(-i))=0$ (where $d$ is the dimension of $G$) by Nakano vanishing theorem since $K_{G}\otimes\mathcal{O}_{G}(-i)$ is negative ample for $i=0,1$.
@@
We now show that if some $X\in\mathcal{H}^{Y}_{1}$ is $k-$ extendable, then a general $X_{t}\in\mathcal{H}^{Y}_{1}$ element is extendable. Let $X_{k+1}\supset X_{k}\supset...\supset X_{1}=X$ be the chain of extensions. Since $X_{1}$ is a prime polarized Calabi-Yau threefold, $X_{2}$ is an anticanonically embedded Fano fourfold, $X_{3}$ is $1/2-$ anticanonically embedded Fano and in general $X_{i}$ is a $1/(i-1)-$ anticanonically embedded Fano $(i+2)-$ fold. Since such polarized varieties are smooth points of their respective Hilbert schemes, we have a chain $\mathcal{H}_{k+1}\to\mathcal{H}_{k}\to...\to\mathcal{H}_{1}=\mathcal{H}_{Y}^{1}$ of non-empty irreducible components of the Hilbert schemes containing the extensions where the maps are given by taking hyperplane sections that send $X_{k+1}$ to $X_{k}$. We show that the map $\mathcal{H}_{i}\to\mathcal{H}_{i-1}$ is surjective for $12\leq i\leq k+$ 1. Denote by $G=Y_{k+1}\supset Y_{k}\supset...\supset Y_{1}=Y$, the chain of extensions of $Y$ and recall that $X_{i}$ obtained as a general member of a one parameter family of smoothings of ribbons $\widetilde{Y}_{i}\in\mathcal{H}_{i}$ on $Y_{i}$ with conormal bundle $\mathcal{O}_{Y_{i}}(-1)$. By standard deformation theory arguments (for example see [CLM93, Section $5.4$]), it is enought to show that the map $H^{1}(T_{X_{i}})\to H^{1}(T_{X_{i}}|_{X_{i-1}})$ is surjective. Hence it is enough to show that $H^{2}(T_{X_{i}}(-1))=0$. Now $H^{2}(T_{X_{i}}(-1))=H^{i}(\Omega_{X_{i}}(K_{X_{i}}\otimes\mathcal{O}_{X_{i}}(1))=0$ for $i\geq 3$ by Nakano vanishing theorem since $K_{X_{i}}\otimes\mathcal{O}_{X_{i}}(1)=(i-2/i-1)K_{X_{i}}$ and is hence negative ample for $i\geq 3$. For $i=2$, we need to show that $H^{2}(\Omega_{X_{2}})=0$. Now $X_{2}$ is obtained as a smoothing of a ribbon over $Y_{2}$ with conormal bundle $\mathcal{O}_{Y_{2}}(-1)$ since the composition of the map $\varphi$ as below
@@
deforms to an embedding, where $\pi$ is a double cover branched along a smooth member of $|\mathcal{O}_{Y_{2}}(2)|$. Hence $X_{2}$ is a deformation of $W_{2}$ and hence it is enough to show that $H^{2}(\Omega_{W_{2}})=0$. By [Par91], we need to show that $H^{2}(\Omega_{Y_{2}})=H^{2}(\Omega_{Y_{2}}(\textrm{log}(B)\otimes\mathcal{O}_{Y_{2}}(-1))=0$. Since $Y_{2}$ is a Fano fourfold which is a linear section of a Mukai variety, we have by Lefschetz theorem on Hodge numbers, that $h^{1,2}(Y_{2})=0$. The latter vanishing amounts to showing $H^{2}(\Omega_{Y_{2}}\otimes\mathcal{O}_{Y_{2}}(-1)=H^{2}(\mathcal{O}_{Y_{2}}(-1)|_{B})=0$. The former is once again zero due to Nakano vanishing theorem. For the latter we need to show that $H^{2}(\mathcal{O}_{Y_{2}}(-1))=H^{3}(\mathcal{O}_{Y_{2}}(-1)\otimes\mathcal{O}_{Y_{2}}(-B))=0$, both of which follows from Kodaira vanishing theorem. $\blacksquare$
@@
For the rest of the families, i.e, $1.10,1.12,1.13-1.15$, the anticanonical embedding of the corresponding Fano-varieties are not smoothly extendable. But some of them are extendable to singular arithmetically Gorenstein normal varieties (see [CLM98], [CDS20]). To apply the above method, one therefore must have a theory of existence and smoothing of embedded ribbons on singular arithmetically Gorenstein normal varieties.
@@
Applying [KLM11, Remark $3.1$], where the authors apply Zak-Lv́ovsky’s theorem for veronese embeddings induced by multiples of a very ample line bundle, one can deduce the following weaker bounds for $l_{Y}$ as follows:
@@
    We first show that a general Calabi-Yau in $\mathcal{H}_{l}^{Y}$ and hence a general canonical surface section in $\mathcal{S}_{l}^{Y}$ for $l\geq l_{Y}$ is projectively normal. By [GP97, Corollary $1.1$], for a Calabi-Yau threefold $X$ and an ample and base point free line bundle $B$ on $X$, $lB$ is projectively normal for $l\geq 4$. For a general $X_{t}\in\mathcal{H}_{l}^{Y}$, the generator $A_{t}$ of the Picard group is base point free and from the tables 1, 2, we see that $l_{Y}\geq 4$ for all deformation type $Y$ other than $1.4-1.10$. Now once again, by [GP97, Theorem $1$], for a Calabi-Yau threefold $X$ and an ample and base point free line bundle $B$ on $X$, $lB$ is projectively normal for $l\geq 3$, unless the morphism by $|B|$ maps $X$, $2:1$ onto $\mathbb{P}^{3}$. In our case therefore for families $1.4-1.10$, $h^{0}(lA_{t})>4$ by Proposition 3.1, part $(3)$ and hence are projectively normal for $l\geq 3$. By [Pin74, remark $7.6$ (iii)], the cone $C(Z)\subset\mathbb{P}^{N_{l}}$ over a general hyperplane section $Z\subset\mathbb{P}^{N_{l}-1}$ lies in every irreducible component of the Hilbert scheme in $\mathbb{P}^{N}$ that dominates $\mathcal{S}_{l}^{Y}$. But now by Lemma 2.5 the Hilbert point of such a cone is smooth. Hence part $(4)$ now follows.
@@
We prove for $Y$ in family $1.7$. Let $G_{n}$ denote the Grassmannian Gr$(2,n+2)$ so that we have Gr $(2,6)$ is denoted by $G_{4}$. $G_{4}$ is a Fano variety of dimension $8$ and index $6$ (see for ex [Man15], Section $(2)$) while its Picard group is generated by the line bundle $\mathcal{O}_{G_{4}}(1)$ corresponding to the Plucker embedding. Hence $K_{G_{4}}=\mathcal{O}_{G_{4}}(-6)$. On the other hand $Y$ is the intersection of $G_{4}$ in its Plucker embedding by a codimension $5$ subspace. Therefore, $K_{Y}=\mathcal{O}_{Y}(-1)$ where $\mathcal{O}_{Y}(1)$ is the restriction of $\mathcal{O}_{G_{4}}(1)$ to $Y$. Let us show $H^{i}(T_{Y}(-l))=0$ for $l\geq 2$. Noting that the normal bundle of $Y$ inside $G_{4}$ is $\mathcal{O}_{Y}(1)^{\oplus 5}$ we have the exact sequence
@@
We have that $H^{i}(\mathcal{Q}^{*}(1-l))=H^{6-i}(\mathcal{Q}(l-6))=H^{6-i}(\Sigma^{(l-6,l-6)}\mathcal{U}^{*}\otimes\Sigma^{(1,0,0)}\mathcal{Q})=H^{6-i}(L_{(l-6,l-6,1,0,0)})$. Setting $\alpha=(l-6,l-6,1,0,0)$ and $\rho=(5,4,3,2,1)$, we have that $\alpha+\rho=(l-1,l-2,4,2,1)$. For any $2\leq l\leq 6$, the entries of $\alpha+\rho$ are not distinct while if $l\geq 7$, the entries are in strictly decreasing order and consequently the permutation that arranges the sequence in strictly decreasing order is identity whose length is zero. Hence $H^{6-i}(L_{(l-6,l-6,1,0,0)})=0$ by the Borel-Weil-Bott theorem (see [Kuz05, Section $2.6$]). Now consider the Koszul resolution of the ideal sheaf $\mathcal{I}_{Y}$
@@
(see https://mathoverflow.net/questions/300999/kozsul-resolution-of-mathcalo-x)
@@
We have that for $l=4$, $h^{0}(\mathcal{O}_{Y}(6-l))=h^{0}(\mathcal{O}_{Y}(2))=11$, for $l=5$, $h^{0}(\mathcal{O}_{Y}(6-l))=h^{0}(\mathcal{O}_{Y}(1))=3$, for $l=6$, $h^{0}(\mathcal{O}_{Y}(6-l))=h^{0}(\mathcal{O}_{Y})=1$, while if $l\geq 7$, $h^{0}(\mathcal{O}_{Y}(6-l))=0$. Further for any value of $l$, $h^{1}(\mathcal{O}_{Y}(6-l))=0$. Now note that if $\overline{\Omega}_{\mathbb{P}}$ denotes the sheaf of regular differential forms as defined in [Dol82, Section $(2)$], then $T_{\mathbb{P}}=\overline{\Omega}_{\mathbb{P}}^{*}$. To see this, further note that if $i:\mathbb{P}_{\textrm{sm}}\hookrightarrow\mathbb{P}$ denotes the inclusion of smooth locus, then $i_{*}(T_{\mathbb{P}_{\textrm{sm}}})=T_{\mathbb{P}}$ since both are reflexive sheaves that agree on the smooth locus which has codimension at least $2$. Further it is shown in [Dol82, $2.2.4$], that $i_{*}(\Omega_{\mathbb{P}_{\textrm{sm}}})=\overline{\Omega_{\mathbb{P}}}$. We have $i_{*}(\Omega_{\mathbb{P}_{\textrm{sm}}}^{*})=i_{*}(\Omega_{\mathbb{P}_{\textrm{sm}}})^{*}$ because both are reflexive sheaves that agree on on the smooth locus which has codimension at least $2$. So finally $T_{\mathbb{P}}=i_{*}(T_{\mathbb{P}_{\textrm{sm}}})=i_{*}(\Omega_{\mathbb{P}_{\textrm{sm}}}^{*})=i_{*}(\Omega_{\mathbb{P}_{\textrm{sm}}})^{*}=\overline{\Omega}_{\mathbb{P}}^{*}$. Now we dualize the Euler sequence
@@
Now restricting to $Y$, tensoring by $\mathcal{O}_{Y}(-l)$ and taking cohomology we have by [Dol82, $1.4$], that $H^{1}(T_{\mathbb{P}}(-l)|_{Y})=0$ for any $l$. Further using results on surjectivity of multiplication maps of global sections of ample line bundles on toric varieties (and hence on weighted projective spaces) ( [NO02]), we have that $H^{2}(T_{\mathbb{P}}(-l)|_{Y})=0$ for $l\geq 6$. Hence our result follows. $\blacksquare$

### v2

@@ 3. Calabi-Yau ribbons on Fano threefolds and deformations of Calabi-Yau double covers of Fano threefolds
Since $H^{2}(\mathcal{O}_{X})=0$, we have that the map $H^{1}(\mathcal{E}_{L})\to H^{1}(T_{X})$ between the first order deformations of the polarized pair $(X,lA)$ and the first order deformations of $X$ is surjective while the respective maps of obstruction spaces is injective. So we have that the forgetful map between the functor of deformations of the polarized pair to the functor of deformations of $X$ is smooth. By the $T^{1}-$ lifting criterion by [38], $X$ is unobstructed and hence $(X,lA)$ is unobstructed as well. So the pair $(X,lA)$ is contained in a unique irreducible component of the moduli space of smooth polarized Calabi-Yau threefolds. By the condition of the $(4)$ both of the pairs $(X_{t},\mathcal{O}_{X_{t}}(1))$ and $(X^{\prime}_{s},\mathcal{O}_{X^{\prime}_{s}}(1))$ are also contained the same component of smooth polarized Calabi-Yau threefolds. So there is a one parameter family $(\mathcal{X},\mathcal{O}_{\mathcal{X}}(1))$ where $f:\mathcal{X}\to T$ is a scheme over $T$, such that $(\mathcal{X}_{t},\mathcal{O}_{\mathcal{X}_{t}}(1))=(X_{t},\mathcal{O}_{X_{t}}(1))$ while $(\mathcal{X}_{s},\mathcal{O}_{\mathcal{X}_{s}}(1))=(X^{\prime}_{s},\mathcal{O}_{X^{\prime}_{s}}(1))$. But now since $H^{i}(\mathcal{O}_{\mathcal{X}_{p}}(1))$ for $i>0$ for any $p\in T$, we have that $h^{0}(\mathcal{O}_{\mathcal{X}_{p}}(1))$ is constant for $p\in T$. Therefore $f_{*}(\mathcal{O}_{\mathcal{X}}(1))$ is locally free of rank $N_{l}+1$ and $\mathcal{X}\to\mathbb{P}(f_{*}(\mathcal{O}_{\mathcal{X}}(1)))$ is a $T-$ morphism for $T$ irreducible such that at $t\in T$ we get the embedding $X_{t}\subset\mathbb{P}^{N_{l}}$ while at $s\in T$, we get the embedding $X^{\prime}_{s}\subset\mathbb{P}^{N_{l}}$. Therefore the $X_{t}\subset\mathbb{P}^{N_{l}}$ and $X^{\prime}_{s}\subset\mathbb{P}^{N_{l}}$ belong to the same irreducible component of the Hilbert scheme. $\blacksquare$
@@ 4. Applications to extendability of Calabi-Yau threefolds and canonical surfaces
Proof. In the case, $1.2-1.4$, the general elements of $\mathcal{H}^{Y}_{1}$ are complete intersections and hence the result follows. We concentrate on $1.5-1.11$. We first show that in each case, a special locus in $\mathcal{H}^{Y}_{j}$ is $k-$ extendable for $k$ as mentioned in the table. In each of the cases the $k-$ th step of the respective embedding $Y\xhookrightarrow{|\mathcal{O}_{Y}(j)|}\mathbb{P}^{M}$ is a Fano variety $G\hookrightarrow\mathbb{P}^{M+k}$ as follows (see [32], [33] or [9] or [4]).
@@
This implies that $T_{\mathbb{P}^{M+k}|_{G}}(-1)$ is base point free and hence $N_{G/\mathbb{P}^{M+k}}(-1)$ is base point free. Further the rank of $N_{G/\mathbb{P}^{M}}(-1)$ which is $M-3$ is greater than $3$. Hence there exists a nowhere vanishing section of $H^{0}(N_{G/\mathbb{P}^{M+k}}(-1))$ which corresponds to an embedded ribbon $\widetilde{G}\subset\mathbb{P}^{M+k}$ on $G\subset\mathbb{P}^{M+k}$. Hence we get a very ample polarization $\mathcal{O}_{\widetilde{G}}(1)$ which restricts to $\mathcal{O}_{G}(1)$ on $G$. Note that $H^{0}(\mathcal{O}_{\widetilde{G}}(1))=H^{0}(\mathcal{O}_{G}(1))+1$ and hence the embedding $\widetilde{G}\subset\mathbb{P}^{M+k}$ is actually induced by a sublinear series of $\mathcal{O}_{\widetilde{G}}(1)$ of codimension $1$. Now embed $\widetilde{G}\subset\mathbb{P}^{M+k+1}$ by the complete linear series of $\mathcal{O}_{\widetilde{G}}(1)$. To show that $\widetilde{G}$ has an embedded smoothing in $\mathbb{P}^{M+k+1}$, we work as in Proposition 3.1. By [5, Proposition $2.16$] (which uses [19, Proposition $1.4$]) we need to show that $|\mathcal{O}_{G}(2)|$ has a smooth member and prove the vanishing of $H^{1}(\mathcal{O}_{G}(2)),H^{1}(N_{G/\mathbb{P}^{M}}(-1))$ and $H^{1}(N_{G/\mathbb{P}^{M}})$. The first one follows from Kodaira vanishing theorem. From the above two exact sequences, for the other two vanishings, one needs to show the vanishing of $H^{1}(T_{\mathbb{P}^{M}}|_{G}(-i))$ and $H^{2}(T_{G}(-i))$ for $i=0,1$. The first one boils down to the vanishings of $H^{1}(\mathcal{O}_{G}(i))$ for $i=0,1$ and $H^{2}(\mathcal{O}_{G}(-i))$ for $i=0,1$, both of which are true since $G$ is Fano. Finally $H^{2}(T_{G}(-i))=H^{d-2}(\Omega_{G}\otimes K_{G}\otimes\mathcal{O}_{G}(-i))=0$ (where $d$ is the dimension of $G$) by Nakano vanishing theorem since $K_{G}\otimes\mathcal{O}_{G}(-i)$ is negative ample for $i=0,1$.
@@
We now show that if some $X\in\mathcal{H}^{Y}_{1}$ is $k-$ extendable, then a general $X_{t}\in\mathcal{H}^{Y}_{1}$ element is extendable. Let $X_{k+1}\supset X_{k}\supset...\supset X_{1}=X$ be the chain of extensions. Since $X_{1}$ is a prime polarized Calabi-Yau threefold, $X_{2}$ is an anticanonically embedded Fano fourfold, $X_{3}$ is $1/2-$ anticanonically embedded Fano and in general $X_{i}$ is a $1/(i-1)-$ anticanonically embedded Fano $(i+2)-$ fold. Since such polarized varieties are smooth points of their respective Hilbert schemes, we have a chain $\mathcal{H}_{k+1}\to\mathcal{H}_{k}\to...\to\mathcal{H}_{1}=\mathcal{H}_{Y}^{1}$ of non-empty irreducible components of the Hilbert schemes containing the extensions where the maps are given by taking hyperplane sections that send $X_{k+1}$ to $X_{k}$. We show that the map $\mathcal{H}_{i}\to\mathcal{H}_{i-1}$ is surjective for $12\leq i\leq k+$ 1. Denote by $G=Y_{k+1}\supset Y_{k}\supset...\supset Y_{1}=Y$, the chain of extensions of $Y$ and recall that $X_{i}$ obtained as a general member of a one parameter family of smoothings of ribbons $\widetilde{Y}_{i}\in\mathcal{H}_{i}$ on $Y_{i}$ with conormal bundle $\mathcal{O}_{Y_{i}}(-1)$. By standard deformation theory arguments (for example see [7, Section $5.4$]), it is enought to show that the map $H^{1}(T_{X_{i}})\to H^{1}(T_{X_{i}}|_{X_{i-1}})$ is surjective. Hence it is enough to show that $H^{2}(T_{X_{i}}(-1))=0$. Now $H^{2}(T_{X_{i}}(-1))=H^{i}(\Omega_{X_{i}}(K_{X_{i}}\otimes\mathcal{O}_{X_{i}}(1))=0$ for $i\geq 3$ by Nakano vanishing theorem since $K_{X_{i}}\otimes\mathcal{O}_{X_{i}}(1)=(i-2/i-1)K_{X_{i}}$ and is hence negative ample for $i\geq 3$. For $i=2$, we need to show that $H^{2}(\Omega_{X_{2}})=0$. Now $X_{2}$ is obtained as a smoothing of a ribbon over $Y_{2}$ with conormal bundle $\mathcal{O}_{Y_{2}}(-1)$ since the composition of the map $\varphi$ as below
@@
deforms to an embedding, where $\pi$ is a double cover branched along a smooth member of $|\mathcal{O}_{Y_{2}}(2)|$. Hence $X_{2}$ is a deformation of $W_{2}$ and hence it is enough to show that $H^{2}(\Omega_{W_{2}})=0$. By [36], we need to show that $H^{2}(\Omega_{Y_{2}})=H^{2}(\Omega_{Y_{2}}(\textrm{log}(B)\otimes\mathcal{O}_{Y_{2}}(-1))=0$. Since $Y_{2}$ is a Fano fourfold which is a linear section of a Mukai variety, we have by Lefschetz theorem on Hodge numbers, that $h^{1,2}(Y_{2})=0$. The latter vanishing amounts to showing $H^{2}(\Omega_{Y_{2}}\otimes\mathcal{O}_{Y_{2}}(-1)=H^{2}(\mathcal{O}_{Y_{2}}(-1)|_{B})=0$. The former is once again zero due to Nakano vanishing theorem. For the latter we need to show that $H^{2}(\mathcal{O}_{Y_{2}}(-1))=H^{3}(\mathcal{O}_{Y_{2}}(-1)\otimes\mathcal{O}_{Y_{2}}(-B))=0$, both of which follows from Kodaira vanishing theorem. $\blacksquare$
@@
For the rest of the families, i.e, $1.10,1.12,1.13-1.15$, the anticanonical embedding of the corresponding Fano-varieties are not smoothly extendable. But some of them are extendable to singular arithmetically Gorenstein normal varieties (see [9], [6]). To apply the above method, one therefore must have a theory of existence and smoothing of embedded ribbons on singular arithmetically Gorenstein normal varieties.
@@
Applying [25, Remark $3.1$], where the authors apply Zak-Lv́ovsky’s theorem for veronese embeddings induced by multiples of a very ample line bundle, one can deduce the following weaker bounds for $l_{Y}$ as follows:
@@
    We first show that a general Calabi-Yau in $\mathcal{H}_{l}^{Y}$ and hence a general canonical surface section in $\mathcal{S}_{l}^{Y}$ for $l\geq l_{Y}$ is projectively normal. By [22, Corollary $1.1$], for a Calabi-Yau threefold $X$ and an ample and base point free line bundle $B$ on $X$, $lB$ is projectively normal for $l\geq 4$. For a general $X_{t}\in\mathcal{H}_{l}^{Y}$, the generator $A_{t}$ of the Picard group is base point free and from the tables 1, 2, we see that $l_{Y}\geq 4$ for all deformation type $Y$ other than $1.4-1.10$. Now once again, by [22, Theorem $1$], for a Calabi-Yau threefold $X$ and an ample and base point free line bundle $B$ on $X$, $lB$ is projectively normal for $l\geq 3$, unless the morphism by $|B|$ maps $X$, $2:1$ onto $\mathbb{P}^{3}$. In our case therefore for families $1.4-1.10$, $h^{0}(lA_{t})>4$ by Proposition 3.1, part $(3)$ and hence are projectively normal for $l\geq 3$. By [37, remark $7.6$ (iii)], the cone $C(Z)\subset\mathbb{P}^{N_{l}}$ over a general hyperplane section $Z\subset\mathbb{P}^{N_{l}-1}$ lies in every irreducible component of the Hilbert scheme in $\mathbb{P}^{N}$ that dominates $\mathcal{S}_{l}^{Y}$. But now by Lemma 2.5 the Hilbert point of such a cone is smooth. Hence part $(4)$ now follows.
@@
We prove for $Y$ in family $1.7$. Let $G_{n}$ denote the Grassmannian Gr$(2,n+2)$ so that we have Gr $(2,6)$ is denoted by $G_{4}$. $G_{4}$ is a Fano variety of dimension $8$ and index $6$ (see for ex [31], Section $(2)$) while its Picard group is generated by the line bundle $\mathcal{O}_{G_{4}}(1)$ corresponding to the Plucker embedding. Hence $K_{G_{4}}=\mathcal{O}_{G_{4}}(-6)$. On the other hand $Y$ is the intersection of $G_{4}$ in its Plucker embedding by a codimension $5$ subspace. Therefore, $K_{Y}=\mathcal{O}_{Y}(-1)$ where $\mathcal{O}_{Y}(1)$ is the restriction of $\mathcal{O}_{G_{4}}(1)$ to $Y$. Let us show $H^{i}(T_{Y}(-l))=0$ for $l\geq 2$. Noting that the normal bundle of $Y$ inside $G_{4}$ is $\mathcal{O}_{Y}(1)^{\oplus 5}$ we have the exact sequence
@@
We have that $H^{i}(\mathcal{Q}^{*}(1-l))=H^{6-i}(\mathcal{Q}(l-6))=H^{6-i}(\Sigma^{(l-6,l-6)}\mathcal{U}^{*}\otimes\Sigma^{(1,0,0)}\mathcal{Q})=H^{6-i}(L_{(l-6,l-6,1,0,0)})$. Setting $\alpha=(l-6,l-6,1,0,0)$ and $\rho=(5,4,3,2,1)$, we have that $\alpha+\rho=(l-1,l-2,4,2,1)$. For any $2\leq l\leq 6$, the entries of $\alpha+\rho$ are not distinct while if $l\geq 7$, the entries are in strictly decreasing order and consequently the permutation that arranges the sequence in strictly decreasing order is identity whose length is zero. Hence $H^{6-i}(L_{(l-6,l-6,1,0,0)})=0$ by the Borel-Weil-Bott theorem (see [27, Section $2.6$]). Now consider the Koszul resolution of the ideal sheaf $\mathcal{I}_{Y}$
@@
We have that for $l=4$, $h^{0}(\mathcal{O}_{Y}(6-l))=h^{0}(\mathcal{O}_{Y}(2))=11$, for $l=5$, $h^{0}(\mathcal{O}_{Y}(6-l))=h^{0}(\mathcal{O}_{Y}(1))=3$, for $l=6$, $h^{0}(\mathcal{O}_{Y}(6-l))=h^{0}(\mathcal{O}_{Y})=1$, while if $l\geq 7$, $h^{0}(\mathcal{O}_{Y}(6-l))=0$. Further for any value of $l$, $h^{1}(\mathcal{O}_{Y}(6-l))=0$. Now note that if $\overline{\Omega}_{\mathbb{P}}$ denotes the sheaf of regular differential forms as defined in [12, Section $(2)$], then $T_{\mathbb{P}}=\overline{\Omega}_{\mathbb{P}}^{*}$. To see this, further note that if $i:\mathbb{P}_{\textrm{sm}}\hookrightarrow\mathbb{P}$ denotes the inclusion of smooth locus, then $i_{*}(T_{\mathbb{P}_{\textrm{sm}}})=T_{\mathbb{P}}$ since both are reflexive sheaves that agree on the smooth locus which has codimension at least $2$. Further it is shown in [12, $2.2.4$], that $i_{*}(\Omega_{\mathbb{P}_{\textrm{sm}}})=\overline{\Omega_{\mathbb{P}}}$. We have $i_{*}(\Omega_{\mathbb{P}_{\textrm{sm}}}^{*})=i_{*}(\Omega_{\mathbb{P}_{\textrm{sm}}})^{*}$ because both are reflexive sheaves that agree on on the smooth locus which has codimension at least $2$. So finally $T_{\mathbb{P}}=i_{*}(T_{\mathbb{P}_{\textrm{sm}}})=i_{*}(\Omega_{\mathbb{P}_{\textrm{sm}}}^{*})=i_{*}(\Omega_{\mathbb{P}_{\textrm{sm}}})^{*}=\overline{\Omega}_{\mathbb{P}}^{*}$. Now we dualize the Euler sequence
@@
Now restricting to $Y$, tensoring by $\mathcal{O}_{Y}(-l)$ and taking cohomology we have by [12, $1.4$], that $H^{1}(T_{\mathbb{P}}(-l)|_{Y})=0$ for any $l$. Further using results on surjectivity of multiplication maps of global sections of ample line bundles on toric varieties (and hence on weighted projective spaces) ( [35]), we have that $H^{2}(T_{\mathbb{P}}(-l)|_{Y})=0$ for $l\geq 6$. Hence our result follows. $\blacksquare$